A 35?F capacitor is connected across an AC generator that produces a peak voltage of 6.0V . The peak current is 0.40

What is the oscillation frequency in Hz? could you define the variables and break it down please?


A
Answer #1

media%2F46c%2F46cff3ab-2d3f-46d3-b18b-31

Answer question
Answer #2

Capacitive reactance

Xc=V/I =6/0.4

Xc=15 ohms

since

Xc=1/2pifC

=>f=1/2pi*C*Xc =1/2pi*(35*10^-6)*15

f=303.15 Hz
Answer #3

impedance=6/0.4=15 ohms


so 1/(2*pi*f*35*10^(-6))=15


frequency=f=303.15 Hz